OpenStudy (anonymous):

simplify the expression of -44=w+32

OpenStudy (anonymous):

Hey taaaaaay122, could you elaborate a little more on this question?

OpenStudy (anonymous):

W= -76

OpenStudy (satish27may):

Given: -44 = w+32 w=-44-32 w=76